A sphere is inscribed in a cube of length 5cm. Find the difference between the volume of the cube and the sphere.
Answer by stanbon(57352)

A sphere is inscribed in a cube of length 5cm. Find the difference between the volume of the cube and the sphere.
------------------------------
Volume of the cube is 5^3 = 125 cm^3
---
radius of the sphere = 5/2 cm
---
Volume of the sphere = (4/3)pi^(5/2)^3
-----
Volume between the sphere and the cube = 125-[(4/3)*pi*(5/2)^3]
= 59.55 cm^3
================
Cheers,
Stan H.